OpenStudy (anonymous):

Which of the following sequences is not arithmetic? 1, 2, 3, 4, ... 2, 4, 6, 8, ... 0, -2, -4, -6, ... 1, 4, 9, 16, ...

OpenStudy (anonymous):

the last one because it's a geometric sequence

OpenStudy (jaredstone4):

So for sequences, you have two types: arithmetic and geometric. The key difference between them is how you get from one term to the next. Arithmetic sequences have a "d", or common difference, and geometric has "r" (common ratio I think?) Anyway difference means subtract, right? So to find d you subtract one term from the one before it. In the first answer choice, 2-1 =1, 3-2 =1, and 4-3 =1. 1 is your d which makes this an arithmetic sequence. Same goes for the second and third choices because if you go through the same process, your d will be 2 and -2 respectively. For the 4th answer choice, that doesn't work. 4-1 =3 9-4 = 5 and 16-9 = 7. This is not a common difference so it is not arithmetic and therefore your answer.